Measures of liquidity, Solvency, and Profitability

The comparative financial statements of Marshall Inc. are as follows. The market price of Marshall common stock was $ 56 on December 31, 20Y2.

Marshall Inc.
Comparative Retained Earnings Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Retained earnings, January 1 $1,793,525 $1,515,775
Net income 425,600 310,400
Total $2,219,125 $1,826,175
Dividends:
On preferred stock $13,300 $13,300
On common stock 19,350 19,350
Total dividends $32,650 $32,650
Retained earnings, December 31 $2,186,475 $1,793,525


Marshall Inc.
Comparative Income Statement
For the Years Ended December 31, 20Y2 and 20Y1
   20Y2    20Y1
Sales $2,205,330 $2,031,930
Cost of goods sold 854,100 785,770
Gross profit $1,351,230 $1,246,160
Selling expenses $407,710 $529,130
Administrative expenses 347,300 310,760
Total operating expenses $755,010 $839,890
Income from operations $596,220 $406,270
Other revenue 31,380 25,930
$627,600 $432,200
Other expense (interest) 144,000 79,200
Income before income tax $483,600 $353,000
Income tax expense 58,000 42,600
Net income $425,600 $310,400


Marshall Inc.
Comparative Balance Sheet
December 31, 20Y2 and 20Y1
   20Y2    20Y1
Assets
Current assets
Cash $450,950 $432,610
Marketable securities 682,520 716,890
Accounts receivable (net) 430,700 401,500
Inventories 321,200 248,200
Prepaid expenses 85,324 86,520
Total current assets $1,970,694 $1,885,720
Long-term investments 1,262,961 769,607
Property, plant, and equipment (net) 2,160,000 1,944,000
Total assets $5,393,655 $4,599,327
Liabilities
Current liabilities $597,180 $1,005,802
Long-term liabilities:
Mortgage note payable, 8% $810,000 $0
Bonds payable, 8% 990,000 990,000
Total long-term liabilities $1,800,000 $990,000
Total liabilities $2,397,180 $1,995,802
Stockholders' Equity
Preferred $0.70 stock, $20 par $380,000 $380,000
Common stock, $10 par 430,000 430,000
Retained earnings 2,186,475 1,793,525
Total stockholders' equity $2,996,475 $2,603,525
Total liabilities and stockholders' equity $5,393,655 $4,599,327

Required:

Determine the following measures for 20Y2, rounding to one decimal place, except for dollar amounts, which should be rounded to the nearest cent. Use the rounded answer of the requirement for subsequent requirement, if required. Assume 365 days a year.

1. Working capital $
2. Current ratio
3. Quick ratio
4. Accounts receivable turnover
5. Number of days' sales in receivables days
6. Inventory turnover
7. Number of days' sales in inventory days
8. Ratio of fixed assets to long-term liabilities
9. Ratio of liabilities to stockholders' equity
10. Times interest earned
11. Asset turnover
12. Return on total assets %
13. Return on stockholders’ equity %
14. Return on common stockholders’ equity %
15. Earnings per share on common stock $
16. Price-earnings ratio
17. Dividends per share of common stock $
18. Dividend yield %
Answer #1

1) Working capital = Current assets-Current liabilities = 1970694-597180 = 1373514

2) Current ratio = Current asset/Current liabilities = 1970694/597180 = 3.3

3) Quick ratio = Quick asset/Current liabilities = 1564170/597180 = 2.62:1

4) Account receivable turnover = 2205330/416100 = 5.3 Times

Note: Please post whole question in the group of 4-4 sub parts as per HOMEWORKLIB RULES